100 Mhz reference unlocks


     Observatory devices use 5, 10, 20, and 100Mhz reference signals to lock themselves to the hydrogen maser. The maser provides 5 Mhz (or is it 1..). This  is converted to 10,20 and 100 Mhz. The 100 Mhz reference is used to drive the correlators (interim and wapp) as well as all of the equipment upstairs in the dome (transmitter and synthesizer for the first lo).

    On 25nov02 peter backus noticed that the seti test tone was drifting. Since the drift rate did not match that of the earth's rotation (and it was monday) he decided it was probably a problem with the LO system rather than ET. On 28nov02 we found that the AO 100 MHz reference was drifting. We temporarily replaced it with a synthesizer locked to the 10 Mhz.

    The 100 MHz reference is generated using a PLL driven from the 5 Mhz. Ganesh pointed out that most VCO's drift with time and all that was needed was to tune the oscillator.  The loop bandwidth  was about 60 Hz. We adjusted the  VCO putting the 5 Mhz at the center of the loop bandwidth.  After this, the system locked up  ok.

    The question is when did the 100 Mhz unlock? A time line of the events are:

20nov02: sband radar ranging to titan with echo --> it's locked
22nov02: vlba holography run. Need to see if they get fringes.
23nov02: sband radar run with narrow echo but 50 Hhz from the ephemeris prediction.
25nov02: peter finds the system unlocked
28nov02: 17:00 replace pll with synthesizer.
02dec02: take data to characterize drift, then retune vco.

        The 23nov02 echo probably means that it was locked on the 23nov02.

    I took some data on 02dec02 before we adjusted the vco. I let the system warm up for about 1 hour before running. A test tone was transmitted from the birdie maker (in the bowl) at 2380.005 Mhz. This was mixed to base band and sampled at 100 Khz bandwidth for 60 seconds. The plots show the wandering of the tone at sband:

    When peter backus looked at the signal on 25nov02 he saw a 22 second period between the sweeping. On 27nov02 we saw a 15 second period between the sweeps. On 02dec02 after an hour warmup, the sweep period was 2.8 seconds. So the unlocked oscillator was definitely changing with time (probably do to temperature).
processing: x101/021202/doit.pro